# Heterogeneous Treatment Effects and

### v3

@@ 2.1 From Theory to Empirical Research Design
Mapping this model onto the empirical research design, we will assume that $\omega\in\{\omega^{\prime},\omega^{\prime\prime}\}$ denotes a binary treatment, where $\omega^{\prime}$ indicates no exposure to the (adverse) shock and $\omega^{\prime\prime}$ denotes exposure to the shock. The first outcome, $y_{1}(\omega)$, measures voters’ expected utility from the incumbent. It is obviously difficult and rare to measure utility directly, though one could, in principle, elicit willingness-to-pay for the incumbent over the challenger. A voter’s expected utility from a vote for the incumbent is given by:
$y_{i1}(\omega)$ $=\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}$ (5)
@@
The second outcome, $y_{2}$, measures each voter’s vote choice for the incumbent. Vote choice is obviously a more standard outcome in literature on voter behavior. This outcome is given by:
$y_{i2}(\omega)$ =$=\begin{cases}1&\text{ if }\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}-\pi^{C}\geq 0\\ 0&\text{ else }\end{cases}$ (6)
Our model can be represented as a directed acyclic graph (DAG), as depicted in Figure 1. This representation clarifies a number of assumptions of the model. First, the mechanism through which the shock affects voter preferences and behavior is voter learning. This is evident because the only path from the shock ($\omega$) to the outcomes $y_{1}(\omega)$ and $y_{2}(\omega)$ passes through the voters’ posterior beliefs about the incumbent’s type. Voter learning is moderated by voters’ prior beliefs about the incumbent, $\pi_{i}(\overline{\theta})$. Throughout this paper, we will indicate causal moderation through the arrow pointing to a path rather than a node, following the suggestion of Weinberg (2007). [^7] We could define an extra node to indicate this interaction. We do not do so in the interest of parsimony. Our choice reflects the fact that representation of “interaction” effects in DAGs is not standardized (Nilsson et al., 2021). Because our theoretical model encodes more structure (e.g., excluding the possibility of certain causal moderation effects), this notational convention encodes more structure in graphical representations than is standard. For the purposes of exposition, the mediator, posterior beliefs, is unobserved to researchers. This feature is standard in most applied research on adverse shocks and voting behavior and renders explicit mediation analysis infeasible. We will assume, however, that researchers have a measure of $\pi_{i}$ and $v_{i}$ from a baseline survey.
[…]
Figure 1: Directed acyclic graph representation of the model of voter preferences and behavior. The arrow from the prior to the path between the shock and the posterior indicates that the prior moderates voter updating on the incumbent, as in (3).
Empirical researchers typically will not fully understand the causal structure represented in Figure 1. If this were the case, a researcher may mistakenly think that a shock affects assessments of valence, such that the baseline measure of $v_{i}$ is a moderator. (For reference, we depict this alternative DAG that is inconsistent with our model in Figure A.1.) In the absence of measured mediators, the researcher could assess the mechanisms by evaluating differences in conditional (or subgroup) treatment effects. Specifically, we will follow common practice by supposing that researchers estimate conditional average treatment effects (CATEs) at different levels of a moderator, $X$, as follows:
$CATE(y,X)=E[y(\omega=\omega^{\prime\prime})-y(\omega=\omega^{\prime})|X=x]$ (7)
We will say that treatment effects are heterogeneous if for some $x,x^{\prime}\in X$ where $x\neq x^{\prime}$, $CATE(y,X=x)-CATE(y,X=x^{\prime})\neq 0$. When used to detect mechanisms, researchers typically assert that this form of heterogeneity gives evidence of mechanism activation or presence. Indeed, this was the structure of the claims about HTEs and mechanisms in both empirical examples.
@@
Under this model of voter updating and behavior, do heterogeneous treatment effects provide evidence that the relevant mechanism is voter learning? Ex-ante, the empiricists do not know that learning is the active (or operative) mechanism. To learn about mechanisms, many researchers will estimate HTEs, typically pointing to heterogeneity as evidence of mechanism activation. When is this approach valid? When does it yield invalid substantive inferences about mechanisms? To develop intuitions, we evaluate four HTE combinations using moderators $\textbf{X}=\{\Pi,V\}$, where $\Pi$ is the set of all possible values of $\pi_{i}$ and $V$ is the set of all possible values of $v_{i}$, and outcomes $y\in\{y_{1},y_{2}\}$. Remark 1 shows that for the outcome measuring a voter’s expected utility from a vote for the incumbent ($y_{1})$, heterogeneity in CATEs correctly provides evidence that the mechanism is voter learning about the incumbent’s type.
@@
(a) Given $\pi\neq\pi^{\prime}\in\Pi$, $CATE(y_{1},X=\pi)-CATE(y_{1},X=\pi^{\prime})\neq 0$.
(b) Given $v\neq v^{\prime}\in V$, $CATE(y_{1},X=v)-CATE(y_{1},X=v^{\prime})=0$.
(c) If $CATE(y_{1},X=x)-CATE(y_{1},X^{\prime}=x^{\prime})\neq 0$, then $x,x^{\prime}\in\Pi$.
@@
Researchers will detect heterogeneity with moderator $\pi$ (by a). This occurs because the effect of the shock on a voter’s posterior belief is moderated by her prior in our model. This is a straightforwad implication of Bayesian updating. Moreover, they will only detect heterogeneity in $\pi$, not in $v$, (by c) supporting the inference that the learning mechanism produces the observed ATE. In contrast, HTEs are not observed for the (non)-moderator $v$ (by b) because valence and posterior beliefs (the mechanism) are additively separable in voters’ utility. Here, researchers are unlikely to mis-attribute the mechanism through analysis of HTEs.
However, for the vote choice for the incumbent, $y_{2}$, the results from Remark 1 change. First, researchers may observe HTEs for different levels of valence, $v$, as well as for different levels of prior beliefs, $\pi$. This would lead most researchers to infer that the effect of the shock does work through some channel involving valence in addition to a channel involving voter learning.
@@
(a) Given $\pi\neq\pi^{\prime}\in\Pi$, $CATE(y_{2},\pi)-CATE(y_{2},\pi^{\prime})\neq 0$ almost everywhere.
(b) Given $v\neq v^{\prime}\in V$, $CATE(y_{1},v)-CATE(y_{1},v^{\prime})\neq 0$ almost everywhere if $\min\{v,v^{\prime}\}<\pi^{C}$.
(c) If $CATE(y_{1},x)-CATE(y_{1},x^{\prime})\neq 0$, then $x,x^{\prime}\in\Pi$ or $x,x^{\prime}\in V$.
Why do we see HTEs in $v$ for vote choice? Recall that each voter votes for the incumbent if and only if $\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}-\pi^{C}\geq 0$. But this binary choice means that voter beliefs and valence are no longer additively separable with respect to vote choice. Consequently, with a sufficiently large sample size (and thus sufficient statistical power), researchers are apt to detect HTEs even when a mechanism is not active. The standard interpretation of this test would lead to a Type-I error in our inference about the activation of the valence mechanism.
@@
    The use of HTEs to measure mechanism activation relies on assumptions about the relationship between moderators and mediators of interest which are typically implicit.
@@
    The contrast between Remarks 1 and 2 in which the theory (and thus mechanism) is fixed but outcomes differ shows that the use of HTEs for assessing mechanism activation depends on the data-generating process behind the outcome of interest.
@@ 3 Framework
We introduce a framework that allows us to analyze the relationship between HTEs—as estimated by treatment-covariate interactions—and the detection of causal mechanisms. Our framework is built upon the potential outcomes framework or Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974). We denote treatment by $Z$. In order to consider HTEs with respect to pre-treatment moderators, denote the vector of measured pre-treatment covariates by $X=(X_{1},X_{2},...,X_{K})\in\mathbb{R}^{K}$. As in our running example, it need not be the case that all $K$ covariates moderate a given treatment effect: in that example, the prior ($\pi_{i}$) moderates the effect but valence ($v_{i}$) does not.
@@
A valid mediator, or mechanism representation, should: (1) be affected by treatment, $Z$, and (2) have a non-zero effect on the outcome. Importantly, the effect of treatment on a mediator or the mediator’s effect on the outcome could vary with some covariate(s), $X_{k}$. We define a mediator as a function denoted by $M(Z;X)$. [^8] Note that potential outcomes are often written as a function of only the manipulated treatment, e.g., $M(Z)$ and $Y(Z)$, to reflect “no causation without manipulation” (Holland, 1986: p. 959). We choose to denote covariates $X$ as arguments to both potential outcomes in order to clarify the structure of covariates and mediators in our potential outcomes. It represents the potential outcomes of causal mediator given treatment $Z$ and covariates $X$. Finally, we denote a potential outcome by $Y(Z,M;X)$.
3.1 Causal Effects
In order to understand when HTEs permit detection of mechanisms, we formalize existing informal conventions. To do so, we decompose the total effect (on a unit, $i$) into direct and indirect effects, as is standard in the causal mediation literature (Imai and Yamamoto, 2013). Let $M$ be the mediator of interest. We analyze the more general case of multiple mmediators in Appendix C. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
$TE(z,z^{\prime};X)$ $=Y(z,M(z;X);X)-Y(z^{\prime},M(z^{\prime};X);X)$ (8)
Treatment effects consist of direct ($DE$) and indirect ($IE$) effects, as follows:  [^9] See Acharya, Blackwell, and Sen (2016) for the identification of the controlled direct effect.
$DE(z,z^{\prime};X)=Y(z,M(z;X);X)-Y(z^{\prime},M(z;X);X)$ (9)
$IE(z,z^{\prime};X)=Y(z^{\prime},M(z;X);X)-Y(z^{\prime},M(z^{\prime};X);X)$ (10)
@@ 3.1 Causal Effects
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediator $M$ at potential outcomes $M(z;X)$. Our preferred interpretation of the direct effect is the composite effect of all other mechanisms (aside from $M$). However, the direct effect could also include unmediated effects of treatment on an outcome. In our running example, there is only one mechanism—voter learning—so the direct effect (all other mechanisms) is zero. This is evident because the treatment has zero effect if we fix the posterior belief to the same level. The indirect effect measures the effect on the outcome that operates by changing the potential outcome of mediator $M$. In our example, the indirect effect measures the effect that passes through the learning mechanism. With a single mechanism, the indirect effect is equivalent to the total effect of treatment.
@@
As is standard, we can re-write the total effect as follows: [^10] Here, we assume that the direct and indirect effects do not vary at different levels of Z. See more discussion by Imai et al. (2011). All results in the paper hold with other decompositions. See additional discussions of general cases in the Appendix C and correlated mechanisms in the Appendix E.
$TE(z,z^{\prime};X)$ $=DE(z,z^{\prime};X)+IE(z,z^{\prime};X),$ (11)
@@
$ATE(z,z^{\prime})$ $=E_{X}[Y(z)-Y(z^{\prime})]$ (12)
$=E_{X}[DE(z,z^{\prime};X)+IE(z,z^{\prime};X)]$ (13)
$:=ADE(z,z^{\prime})+AIE(z,z^{\prime})$ (14)

### v4

@@ 1.1 From Theory to Empirical Research Design
Mapping this model onto the empirical research design, we consider two measures of voting outcomes. The first outcome, $y_{i1}$, measures voters’ (expected) utility from the incumbent. It is obviously difficult and rare to measure utility from a candidate directly. However, one could, in principle, ask a voter to evaluate their incumbent on a 0-100 scale (or elicit willingness-to-pay for the incumbent’s reelection). As in the model, the voter’s expected utility from a vote for the incumbent is given by:
$y_{i1}(c)$ $=-\lambda_{i}c_{i}+a_{i}+\varepsilon_{i}$ (1)
@@
The second outcome, $y_{i2}$, measures each voter’s (self-reported) vote choice for the incumbent. Vote choice is a commonly measured outcome in literature on voter behavior. This outcome is given by:
$y_{i2}(c)$ $=\begin{cases}1&\text{ if }-\lambda_{i}c_{i}+a_{i}+\varepsilon_{i}\geq 0\\ 0&\text{ else }\end{cases}$ (2)
The data generating process of the model is depicted in Figure A.1. As we report in Table 1, empirical researchers often turn to estimation of heterogeneous treatment effects with respect to a pre-treatment covariate to assess the activation of a mechanism. In the context of an experiment, heterogeneity is assessed through the estimation of conditional average treatment effects (CATEs) at different levels of a pre-treatment covariate, $X_{k}$. Given outcomes $j\in\{1,2\}$:
$CATE(y_{j},X_{k})=E[y_{ij}(c_{i}=1)-y_{ij}(c_{i}=0)|X_{ik}=x]$ (3)
We will say that treatment effects are heterogeneous if for some $x,x^{\prime}\in X_{k}$ where $x\neq x^{\prime}$, $CATE(y_{j},X_{k}=x)-CATE(y_{j},X_{k}=x^{\prime})\neq 0$. Importantly, whether treatment effects are homogeneous or heterogeneous in a given covariate, is fundamentally a qualitative classification. This means that even though researchers are using a quantitative metric to evaluate mechanisms, they are fundamentally making a qualitative inference about mechanism activation.
@@
Recall that standard interpretations in the empirical literature use the presence of HTEs as evidence that a mechanism is active and the absence of HTEs to assert that a mechanism is inert. Given our model of voter behavior, we ask: do heterogeneous treatment effects provide evidence that the relevant mechanism is indeed voter distaste for corruption? To develop intuitions, we evaluate four HTE combinations using moderators $X_{1}=\Lambda$ and $X_{2}=A$, where $\Lambda$ is the set of all possible values of $\lambda_{i}$ (corruption aversion) and $A$ is the set of all possible values of $a_{i}$ (partisan alignment with the incumbent), and outcomes $y\in\{y_{1},y_{2}\}$. Remark 1 shows that for the outcome measuring a voter’s expected utility from a vote for the incumbent ($y_{1})$, heterogeneity in CATEs correctly provides evidence that the mechanism is voter distaste for corruption, not some type of corruption-induced partisan-realignment or biased learning that depends on partisanship (e.g., motivated reasoning in Little et al., 2022) (for example) that are not present in our model.
@@
(a) Given $\lambda>\lambda^{\prime}\in\Lambda$, $|CATE(y_{1},X_{1}=\lambda)|>|CATE(y_{1},X_{1}=\lambda^{\prime})|$.
(b) Given $a\neq a^{\prime}\in A$, $CATE(y_{1},X_{2}=a)-CATE(y_{1},X_{2}=a^{\prime})=0$.
(c) If $CATE(y_{1},X_{k}=x)-CATE(y_{1},X_{k}=x^{\prime})\neq 0$, then $X_{k}=X_{1}=\Lambda$.
@@
Researchers will detect heterogeneity with the corruption aversion moderator $\lambda$ (by (a)). Voters whose corruption aversion $\lambda$ is larger value the incumbent less (upon receiving the information) than voters with lower corruption aversion. HTEs are not observed for the partisan alignment (non)-moderator $a$ (by (b)) because ideological alignment (not the mechanism) and distaste for corruption (the mechanism) are additively separable in voters’ utility (in (1)). Here, (c) shows that researchers are unlikely to mis-attribute the mechanism through conventional interpretation of HTEs.
However, for the vote choice for the incumbent, $y_{2}$, the results from Remark 1 change. First, researchers may observe HTEs for different levels of partisan alignment, $a$, as well as for different levels of corruption aversion, $\lambda$. A naïve interpretation might suggest that the effect of the corruption information does work through some channel involving ideological re-alignment or biased learning in addition to a channel involving distaste for corruption.
@@
(a) For $\lambda>\lambda^{\prime}\in\Lambda$, $|CATE(y_{2},\lambda)|>|CATE(y_{2},\lambda^{\prime})|$.
(b) For $a\in A$, $|CATE(y_{2},a=-1)|<|CATE(y_{2},a=1)|<|CATE(y_{2},a=0)|$.
(c) If $CATE(y_{2},X_{k}=x)-CATE(y_{2},X_{k}=x^{\prime})\neq 0$, then $X_{k}=X_{1}=\Lambda$ or $X_{k}=X_{2}=A$.
Voters with stronger aversion to corruption exhibit larger treatment effects on their voting behavior (by (a)). Interestingly, even though distaste for corruption is the unique mechanism (in the model) through which information affects vote choice, we now observe heterogeneous treatment effects across voters of different partisan alignments. Specifically, the treatment effect is largest among neutral/independent voters ($a=0$) voters, while strongly aligned partisans (a=1,1a=-1,1) exhibit smaller changes in voting behavior, by (b). Moreover, the effect is greater for incumbent-aligned voters than for challenger-aligned voters. This pattern is driven by a measurement concern: ceiling and floor effects that emerge when utility is transformed into vote choice. Aligned voters are substantially more likely to vote for or against the incumbent regardless of new information, while moderates hover around the decision threshold, making them more responsive to changes in observed corruption.
Importantly, the smaller swings observed among strong partisans are not due to the presence of an additional mechanism. Rather, it is because their prior disposition already places them near the upper or lower bounds of the probability scale in terms of propensity to vote for the incumbent. Researchers who are unaware of the underlying mechanism might mistakenly conclude that partisan voters are “less responsive" to treatment and, as a result, wrongly infer that they are engaged in biased learing/insensitive to information or that they are less prone to partisan realignment on the basis of corruption information.
More concretely, we see HTEs in partisan affiliation ($a$) for vote choice because voters make a binary choice between the incumbent and challenger. But this binary choice means that the distaste for corruption mechanism and the partisan alignment predictor (which does not moderate the mechanism) are no longer additively separable with respect to vote choice. Consequently, researchers are apt to detect HTEs in partisan alignment even when it does not moderate any mechanism. A naïve interpretation of this test would lead to a Type-I error in our inference about the activation of a mechanism involving partisan alignment.
While Remarks 1 and 2 rely on theoretical analysis of a model, one might ask whether an empirical researcher would be likely to detect this heterogeneity in their data. We therefore conduct a Monte Carlo simulation in Appendix C.3 that uses our theoretical model to guide the data generation process from a hypothetical experiment. We vary: (1) (average) voter support for the incumbent in the electorate ($\overline{y_{2}(0)}\in\{0.1,0.2,...,0.9\}$) by varying the mean of the valence distribution; and (2) sample size in the hypothetical experiment ($n\in\{100,500,1,000\})$. Consistent with our theoretical results, we observe HTEs in the corruption aversion covariate for both outcomes (voter utility and vote choice) and HTEs in partisan affiliation only for the vote choice outcome. More importantly, with the finite sample sizes that we simulate, we show that for high and low values of voter support for the incumbent in the electorate, we are more likely (better powered) to detect HTEs in partisan affiliation than in corruption aversion for the vote choice outcome. This finding holds across multiple estimators of the difference in CATEs (treating $a_{i}$ linearly or as a factor variable). Thus, we cannot solely rely on the statistical properties of our designs to help us screen heterogeneity that is informative about mechanism activation from heterogeneity that is not.
@@
    The use of HTEs to measure mechanism activation relies on assumptions about the relationship between moderators and mechanisms of interest which are typically implicit.
@@
    The contrast between Remarks 1 and 2 in which the theory (and thus mechanism) is fixed but outcomes differ shows that the use of HTEs for assessing mechanism activation depends on the measurement of outcomes of interest.
@@ 2 Framework
2.1 Defining HTEs
Our framework is built upon the potential outcomes framework or Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974). We denote a randomly assigned treatment by $Z\in\{z,z^{\prime}\}$. [^5] All results hold for observational studies in which treatment is conditionally independent of potential outcomes, given a conditioning set of covariates, $X$. In order to consider HTEs with respect to pre-treatment moderators, denote the vector of pre-treatment covariates by $X=(X_{1},X_{2},...,X_{K})\in\mathbb{R}^{K}$. Some of $X_{k}$ are measured. To improve readability and emphasize the key elements of the framework, we omit subscript $i$ from the notation throughout the framewrok and results sections.
@@ 2.1 Defining HTEs
A valid mediator, or mechanism representation, should: (1) be affected by treatment, $Z$, and (2) have a non-zero effect on the outcome. Further, the effect of treatment on a mediator or the mediator’s effect on the outcome could vary with some covariate(s), $X_{k}$. We define a mediator as a function denoted by M(Z;X)M(Z;X). [^6] Note that potential outcomes are often written as a function of only the manipulated treatment, e.g., $M(Z)$ and $Y(Z)$, to reflect “no causation without manipulation” (Holland, 1986: p. 959). We choose to denote covariates $X$ as arguments to both potential outcomes in order to clarify the structure of covariates and mediators in our potential outcomes. It represents the potential outcomes of causal mediator given treatment $Z$ and covariates $X$. While we employ mediators to clarify the link between HTEs and mediation, these mediators (or mechanisms) do not need to be measured in order to analyze HTEs. Finally, we denote a potential outcome by $Y(Z,M;X)$.
We consider the practice of quantifying heterogeneous treatment effects by estimating and comparing CATEs, as documented in Table 1. Following this convention, we consider HTEs with respect to pre-treatment moderators, i.e., for some variable $X_{k}$, where $k\in\{1,2,...,K\}$.
Definition 1 (Conditional Average Treatment Effect).
Consider pre-treatment covariate $X_{k}$. Given that $z\neq z^{\prime}\in Z$, the conditional average treatment effect (CATE) of $Z$ on $Y$ when $X_{k}=x$ is:
$CATE^{Y}(X_{k}=x)=E_{X_{\neg k}}[Y(Z=z,M;X)-Y(Z=z^{\prime},M;X)|X_{k}=x]$
Note that the CATE of a treatment $Z$ is defined with respect to (potential) outcome variable $Y$. We will index estimands by the outcome variable of interest throughout. Given this definition, there exist HTEs when CATEs of $Z$ on $Y$ differ at different values of a covariate $X_{k}$:
Definition 2 (Heterogeneous treatment effects).
HTEs on $Y$ exist with respect to pre-treatment covariate $X_{k}$ if $CATE^{Y}(X_{k}=x)\neq CATE^{Y}(X_{k}=x^{\prime})$ for some $x\neq x^{\prime}\in X_{k}$. [^7] More precisely, the probability measure for the set that contains such $x$ and $x^{\prime}$ is non-zero.
Definitions 1-2 formalize the current practice of comparing CATEs to evaluate whether treatment effects exhibit heterogeneity. However, while our potential outcomes implicitly depend on a mediator, $M$, there is not yet a link to an underlying mechanism.
2.2 Causal Mediation and Indirect Effects
We now develop the mapping between analysis of treatment effect heterogeneity and causal mediation, which seeks to quantify the effect of one or more mechanisms. Specifically, consider a decomposition of the total effect (on a unit, $i$) into direct and indirect effects (Imai and Yamamoto, 2013). Suppose that there exist two mediators (or mechanisms), indexed by $M_{1},M_{2}$. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
$TE^{Y}(z,z^{\prime};X)$ $=Y(z,M_{1}(z;X),M_{2}(z;X);X)-Y(z^{\prime},M_{1}(z^{\prime};X),M_{2}(z^{\prime};X);X)$ (4)
Our notation varies slightly from conventional presentations of mediation that only consider one mechanism (mediator) (Imai et al., 2010a). In the main text, we describe the case with two mechanisms since it is straightforward to generalize this to the special case of one mechanism or to a setting with more than two mechanisms. Further, as above, we continue to index causal effects by the outcome variable, here $Y$. Treatment effects on $Y$ may consist of direct ($DE^{Y}$) and indirect ($IE^{Y}_{1}$ and $IE^{Y}_{2}$) effects, as follows:  [^8] See Acharya et al. (2016) for the identification of the controlled direct effect.
$DE^{Y}(z,z^{\prime};X)=Y(z,M_{1}(z;X),M_{2}(z;X);X)-Y(z^{\prime},M_{1}(z;X),M_{2}(z;X);X)$ (5)
$IE^{Y}_{1}(z,z^{\prime};X)=Y(z^{\prime},M_{1}(z;X),M_{2}(z;X);X)-Y(z^{\prime},M_{1}(z^{\prime};X),M_{2}(z;X);X)$ (6)
$IE^{Y}_{2}(z,z^{\prime};X)=Y(z^{\prime},M_{1}(z^{\prime};X),M_{2}(z;X);X)-Y(z^{\prime},M_{1}(z^{\prime};X),M_{2}(z^{\prime};X);X)$ (7)
@@ 2.2 Causal Mediation and Indirect Effects
The direct effect, $DE^{Y}(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding both mediators $M$ at potential outcomes $M(z;X)$. Our preferred interpretation of the direct effect is the composite effect of any other mechanisms (aside from $M_{1}$ and $M_{2}$). However, the direct effect could also include unmediated effects of treatment on an outcome. In our motivating example, there is only one mechanism—voter distaste for observed corruption—so the direct effect (all other mechanisms) is zero. This is evident because the treatment has zero effect if were fix the mediator, $\lambda_{i}c_{i}$, to a given level. The indirect effect of mechanism $j\in\{1,2\}$ measures the effect on the outcome that operates by changing the potential outcome of mediator $M_{j}$. In our example, the indirect effect measures the effect that passes through the distaste mechanism.
@@
As is standard, we can re-write the total effect as follows: [^9] Here, we assume that the direct and indirect effects do not vary at different levels of Z. See more discussion by Imai et al. (2011). All results in the paper hold with other decompositions, albeit with slightly different interpretations. See additional discussions of general cases in the D and correlated mechanisms in E.
$TE^{Y}(z,z^{\prime};X)$ $=DE^{Y}(z,z^{\prime};X)+IE^{Y}_{1}(z,z^{\prime};X)+IE^{Y}_{2}(z,z^{\prime};X),$ (8)
@@
$ATE^{Y}(z,z^{\prime})$ $=E_{X}[Y(z,M_{1}(z;X),M_{2}(z;X);X)-Y(z^{\prime},M_{1}(z^{\prime};X),M_{2}(z^{\prime};X);X)]$ (9)
$=E_{X}[DE^{Y}(z,z^{\prime};X)+IE^{Y}_{1}(z,z^{\prime};X)+IE^{Y}_{2}(z,z^{\prime};X)]$ (10)
$:=ADE^{Y}(z,z^{\prime})+AIE^{Y}_{1}(z,z^{\prime})+AIE^{Y}_{2}(z,z^{\prime})$ (11)